The sum of n terms of an ap is n squares.find its common difference.

Grade:10

1 Answers

Vikas TU
7 years ago
(n/2)(a + l) = n^2
a + l = 2n
l = 2n – a
l = a + (n-1)d = 2n – a
d = 2(n – a)/(n-1)
where a would be the first term always.